# How to Steer LLM Latents for Hallucination Detection?

### v1

@@ 2 Related Works
Hallucination detection has emerged as a critical area of research, addressing safety concerns of LLMs and their deployment in real-world applications (Huang et al., 2023). A plethora of works address hallucination detection by designing uncertainty scoring functions. For instance, logit-based methods utilize token-level probability as an uncertainty score (Ren et al., 2022; Malinin & Gales, 2021; Kuhn et al., 2023), verbalized methods prompt LLMs to express their uncertainty in human language (Lin et al., 2022b; Xiong et al., 2024), and consistency-based methods assess uncertainty by evaluating the consistency across multiple responses (Manakul et al., 2023; Chen et al., 2024a). Recently, internal state-based methods leverage hidden activations, employing techniques such as contrast-consistent search (Burns et al., 2023) and identifying hallucination subspace (Du et al., 2024). However, these approaches often rely on default LLM embeddings that do not inherently separate truthful and hallucinated data. In contrast, our method aims to shape the latent space through a learnable steering vector for enhanced separation between the two types of data.
@@
Activation engineering enables control over the LLM generation during inference, applying task-specific steering vectors into the model’s internal representation (Zou et al., 2023). For example, several studies mitigate hallucination by shifting activations along the truthful direction identified by analyzing activation differences between contrastive pairs (Li et al., 2024; Chen et al., 2024b; Marks & Tegmark, 2024). Concurrently, representation fine-tuning methods introduce learning task-specific interventions on linear subspaces of hidden representations (Wu et al., 2024) or sparse subsets of attention heads (Yin et al., 2024b).
@@ 3 Problem Setup
Following the practical setup in recent work (Du et al., 2024), we utilize unlabeled LLM generations in the wild, which can be collected in vast quantities through user interactions with LLMs. This data can be freely collectible for any deployed LLM system, yet often contains a mixture of truthful and hallucinated content. Formally,
@@ 4.1 How to learn TSV? Initial training phase
To realize the MLE objective, we need to explicitly model the probability distribution $p(c_{i}\mid\Phi_{\text{final}}(\mathbf{h}^{(l)}_{i}+\lambda\mathbf{v}))$. In particular, we model the last-token embeddings at the final layer using a hyperspherical distribution with the unit norm, where truthful and hallucinated data each form distinct clusters. This modeling aligns with the structure of embeddings typically observed after the RMSNorm layer in practical Transformer models (Dubey et al., 2024; Yang et al., 2024), where the norms of the embeddings are similar but directions can vary (see verification in Appendix E). This can be naturally characterized by the von Mises-Fisher distribution, a classical probability distribution in directional statistics (Mardia & Jupp, 2009), which is analogous to spherical Gaussian distributions for features with unit norms. Under this model, the class conditional probability is given by:
@@ 4.2 How to learn TSV? Augmented training phase
where $\mathbf{1}_{M}\in\mathbb{R}^{M}$ denotes an $M$-dimensional vector of ones, $\mathbf{Q}_{m,c}=\frac{1}{M}q(c|\mathbf{r}^{\mathbf{v}}_{m})$ represents an entry of the matrix $\mathbf{Q}\in\mathbb{R}^{M\times 2}$ for assigned joint pseudo-label probabilities, and $\mathbf{P}_{m,c}=\frac{1}{M}p(c|\mathbf{r}^{\mathbf{v}}_{m})$ denotes an entry of $\mathbf{P}\in\mathbb{R}^{M\times 2}$ for joint probabilities estimated by our model after initial training, where $p(c|\mathbf{r}^{\mathbf{v}}_{m})$ is computed with Eq. 4. The first constraint ensures that for each unlabeled sample, the total probability mass of being assigned to two classes adds up to 1. The second constraint ensures that the number of samples assigned to each class matches the expected class probability distribution $\mathbf{w}\in\mathbb{R}^{2}$. Here, $H(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ is the entropy function, and $\epsilon$ is a hyperparameter controlling the smoothness of the assignment. The entropy regularization term enables the computationally efficient Sinkhorn algorithm (Cuturi, 2013) to solve the problem. The minimizer of Eq. 7 can be expressed as:
@@
The learning process described in Section 4.1 is then repeated using the augmented dataset until convergence. We summarize the full algorithm in Appendix F.
@@ 5.1 Setup
We evaluate our method on four generative question-answering (QA) tasks: three open-domain QA datasets–TruthfulQA (Lin et al., 2022a), TriviaQA (Joshi et al., 2017), and NQ Open (Kwiatkowski et al., 2019); and a domain-specific QA dataset–SciQ (Welbl et al., 2017). For evaluation, 25% of the QA pairs from each dataset are reserved for testing. Consistent with Du et al. (2024), 100 QA pairs are used for validation, while the remaining samples simulate the unlabeled training dataset. We randomly sample $N=32$ pairs from TruthfuQA, and 64 pairs from the other datasets to construct an exemplar set, with $K=128$ used for all experiments. Implementation details are provided in Appendix A.
@@
Evaluation. Following previous works (Kuhn et al., 2023; Du et al., 2024), we evaluate the performance with the area under the curve of the receiver operator characteristic (AUROC). We consider the generation truthful when the similarity score between the generation and the reference answer is larger than a threshold of 0.5. Following Lin et al. (2022a), we utilize BLEURT (Sellam et al., 2020) to measure the similarity. Additionally, we show that our method is robust when evaluated using GPT-4o (Hurst et al., 2024) in Section C.2.
@@ 5.2 Main results
In Table 1, we compare TSV with competitive hallucination detection methods from the literature. TSV demonstrates state-of-the-art performance, significantly outperforming other methods on both the LLaMA-3.1-8b and Qwen-2.5-7b models. We show that unsupervised methods often struggle with inconsistent performance across different models and data distributions as the representations in LLMs are not inherently aligned with the hallucination detection task, making them less reliable for safety-critical applications. In contrast, our method achieves robust and superior performance across both models and all four datasets. In particular, TSV outperforms HaloScope by 13.6% on TruthfulQA with LLaMA-3.1-8b. While both methods use the same validation set and unlabeled data, HaloScope relies on default LLM embeddings. By contrast, our method leverages a small exemplar set and shapes the latent space to better align with the hallucination detection task, enabling significantly improved performance while remaining practical. Our method is also computationally efficient at the inference stage with a complexity of $O(m^{2})$, where $m$ is the number of generated tokens. In contrast, some logit and consistency-based methods require multiple sampling, resulting in a higher complexity of $O(Am^{2})$, where $A$ can be over $10$ in practice. Qualitative results are in Appendix D, and experiments with larger models (LLaMA-3.1-70b & Qwen-2.5-14b) are provided in Section C.1.
@@ 5.3 Ablation studies
Computational efficiency of TSV. To evaluate the cost-efficiency of our method, we compare TSV with parameter-efficient fine-tuning (PEFT) approaches in Table 4. Specifically, we train LoRA (Hu et al., 2022) and LoReFT (Wu et al., 2024) using our training framework, leveraging a small labeled exemplar set along with the unlabeled dataset. Our method achieves superior performance while utilizing 8×8$\times$8 × to 800×800$\times$800 × fewer parameters, demonstrating that TSV can effectively shape representations for the hallucination detection task while significantly reducing computational and annotation costs. Training time is detailed in Appendix G.
@@ 6 Conclusion
Ensuring the reliability of LLM is paramount as they are increasingly integrated into high-stakes applications like healthcare, law, and education. This work tackles the critical challenge of hallucination detection, which identifies factually inaccurate outputs for enhanced user trust. We propose a practical method that minimizes computational and labeling costs while enabling a plug-and-play approach for pre-trained LLMs. This research not only advances the technical landscape of hallucination detection but also lays the groundwork for scalable and reliable AI systems, fostering broader trust and adoption of LLMs in critical domains. Our study does not involve human subjects, complies with all legal and ethical standards, and we do not anticipate any potential harmful consequences resulting from our work. Code will be released for reproducibility.
@@ Acknowledgement
We gratefully acknowledge Maxim Khanov, Shawn Im, and Shrey Modi for their valuable comments on the draft. This work is supported by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS2331669, Office of Naval Research under grant number N00014-23-1-2643, and Philanthropic Fund from SFF.
@@
Appendix A Implementation Details and Hyperparameters
A.1 Implementation details (ours)
Following Kuhn et al. (2023), we generate the most likely answer using beam search with 5 beams. Class prototypes $\boldsymbol{\mu}_{c}$ and TSV $\mathbf{v}$ are randomly initialized, and trained in two stages: 20 epochs using only the exemplar set, followed by an additional 20 epochs after augmentation. Training is performed using the AdamW optimizer (Loshchilov, 2019), with a learning rate of 5e-03 and a batch size of 128. We set steering strength $\lambda$ to 5, the concentration parameter of the vMF distribution $\kappa$ to 10, and the EMA decay rate $\alpha$ to 0.99. The number of iterations in the Sinkhorn algorithm is 3, and the regularization parameter $\epsilon$ is set to 0.05. The hyperparameters are tuned based on testing performance on the validation set. The steering location for each model is detailed in Section A.2. For generating responses, we utilize the following input prompt:
Input prompt for generating responses Prompt:
Answer the question concisely:
Q: {question} A:
A.2 Hyperparameters
Table 5: Steering layer index for LLaMA-3.1-8b and Qwen-2.5-7b.
Model Datasets
TruthfulQA TriviaQA SciQ NQ Open
LLaMA-3.1-8b 9 4 8 9
Qwen-2.5-7b 4 6 11 6
Hyperparameters Search space
Steering MHA component {‘mlp’, ‘attn’, ‘res’}
Steering strength ($\lambda$) {0.1, 0.5, 1, 5, 10}
Optimizer {SGD, Adam, AdamW}
Learning rate {1e-04, 2e-04, 5e-04, 1e-03, 2e-03, 5e-03, 1e-02}
Batch size {32, 64, 128}
Initial training epochs ($n_{\text{initial}}$) {5, 10, 20, 40}
Augmented training epochs ($n_{\text{augmented}}$) {5, 10, 20, 40}
EMA decay rate ($\alpha$) {0, 0.5, 0.9, 0.95, 0.99, 1}
Concentration parameter ($\kappa$) {0.1, 1, 5, 10, 100}
A.3 Implementation details (baselines)
Verbalized Prompt:
Q: {question}
Self-evaluation Prompt:
Q: {question}
Appendix B More Details of the Benchmarks
Appendix C Ablation Studies
C.1 Scalability to larger language models
Table 7: Hallucination detection results on larger LLMs.
Method LLaMA-3.1-70b Qwen-2.5-14b
TruthfulQA SciQ TruthfulQA SciQ
Perplexity 52.3 53.5 64.7 76.0
CCS 63.1 55.5 69.4 80.0
HaloScope 67.4 60.3 74.6 78.9
$\text{SAPLMA}^{\clubsuit}$ 70.2 58.4 83.1 83.8
Ours 76.6 75.0 83.6 89.7
C.2 Evaluation results with GPT-4o
Method LLaMA-3.1-8b Qwen-2.5-7b
TruthfulQA SciQ TruthfulQA SciQ
Perplexity 65.8 41.3 56.3 55.3
CCS 59.5 52.7 58.6 50.3
HaloScope 63.3 66.9 63.4 77.0
$\text{SAPLMA}^{\clubsuit}$ 70.7 72.6 67.7 79.6
TSV (Ours) 75.1 81.6 87.2 85.6
Input prompt for GPT-4o based evaluation Prompt:
C.3 Design choices for the class distribution $\mathbf{w}$
Table 9: Hallucination detection results using different $\mathbf{w}$.
Method LLaMA-3.1-8b Qwen-2.5-7b
TruthfulQA SciQ TruthfulQA SciQ
Uniform 83.2 84.0 87.0 81.2
Estimation 83.7 83.9 87.3 80.6
Oracle 84.3 85.0 87.5 82.6
Ours 84.2 85.8 87.3 82.0
C.4 Results with LLaMA-2-chat-7b
Method TruthfulQA TriviaQA
Perplexity 56.77 72.13
LN-Entropy 61.51 70.91
Semantic Entropy 62.17 73.21
Lexical Similarity 55.69 75.96
EigenScore 51.93 73.98
SelfCKGPT 52.95 73.22
Verbalize 53.04 52.45
Self-evaluation 51.81 55.68
CCS 61.27 60.73
HaloScope 78.64 77.40
TSV (Ours) 80.93 85.20
Appendix D Qualitative Results
Appendix E Embedding Norms
Appendix F Algorithms
F.1 Overall training framework
@@ F.2 Sinkhorn algorithm
Appendix G Compute Resources and Time
G.1 Software and hardware

### v2

# Steer LLM Latents for Hallucination Detection

@@ 2 Related Works
Hallucination detection has emerged as a critical area of research, addressing safety concerns of LLMs and their deployment in real-world applications (Huang et al., 2023). A plethora of works address hallucination detection by designing uncertainty scoring functions. For instance, logit-based methods utilize token-level probability as an uncertainty score (Ren et al., 2022; Malinin & Gales, 2021; Kuhn et al., 2023), verbalized methods prompt LLMs to express their uncertainty in human language (Lin et al., 2022b; Xiong et al., 2024), and consistency-based methods assess uncertainty by evaluating the consistency across multiple responses (Manakul et al., 2023; Chen et al., 2024a). Recently, internal state-based methods such as HaloScope leverage hidden activations to identify hallucination subspace (Du et al., 2024). However, these approaches often rely on default LLM embeddings that do not inherently separate truthful and hallucinated data. In contrast, our method aims to shape the latent space through a learnable steering vector for enhanced separation between the two types of data.
@@
Activation engineering enables control over the LLM generation during inference, applying task-specific steering vectors into the model’s internal representation (Im & Li, 2025). For example, several studies mitigate hallucination by shifting activations along the truthful direction identified by analyzing activation differences between contrastive pairs (Li et al., 2024; Chen et al., 2024b; Marks & Tegmark, 2024). Concurrently, representation fine-tuning methods introduce learning task-specific interventions on linear subspaces of hidden representations (Wu et al., 2024) or sparse subsets of attention heads (Yin et al., 2024b).
@@ 3 Problem Setup
Following the practical setup in recent work HaloScope (Du et al., 2024), we utilize unlabeled LLM generations in the wild, which can be collected in vast quantities through user interactions with LLMs. This data can be freely collected for any deployed LLM system, yet often contains a mixture of truthful and hallucinated content. Formally,
@@ 4.1 How to learn TSV? Initial training phase
To realize the MLE objective, we need to explicitly model the probability distribution $p(c_{i}\mid\Phi_{\text{final}}(\mathbf{h}^{(l)}_{i}+\lambda\mathbf{v}))$. In particular, we model the last-token embeddings at the final layer using a hyperspherical distribution with the unit norm, where truthful and hallucinated data each form distinct clusters. This modeling aligns with the structure of embeddings typically observed after the RMSNorm layer in practical Transformer models (Dubey et al., 2024; Yang et al., 2024), where the norms of the embeddings are similar but directions can vary (see verification in Appendix G). This can be naturally characterized by the von Mises-Fisher distribution, a classical probability distribution in directional statistics (Mardia & Jupp, 2009), which is analogous to spherical Gaussian distributions for features with unit norms. Under this model, the class conditional probability is given by:
@@ 4.2 How to learn TSV? Augmented training phase
where $\mathbf{1}_{M}\in\mathbb{R}^{M}$ denotes an $M$-dimensional vector of ones, $\mathbf{Q}_{m,c}=\frac{1}{M}q(c|\mathbf{r}^{\mathbf{v}}_{m})$ represents an entry of the matrix $\mathbf{Q}\in\mathbb{R}^{M\times 2}$ for assigned joint pseudo-label probabilities, and $\mathbf{P}_{m,c}=\frac{1}{M}p(c|\mathbf{r}^{\mathbf{v}}_{m})$ denotes an entry of $\mathbf{P}\in\mathbb{R}^{M\times 2}$ for joint probabilities estimated by our model after initial training, where $p(c|\mathbf{r}^{\mathbf{v}}_{m})$ is computed with Eq. 4. The first constraint ensures that for each unlabeled sample, the total probability mass of being assigned to two classes adds up to 1. The second constraint ensures that the number of samples assigned to each class matches the expected class probability distribution $\mathbf{w}\in\mathbb{R}^{2}$. Here, $H(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ is the entropy function, and $\epsilon$ is a hyperparameter controlling the smoothness of the assignment, which we set to $0.05$. The entropy regularization term enables the computationally efficient Sinkhorn algorithm (Cuturi, 2013) to solve the problem. The minimizer of Eq. 7 can be expressed as:
@@
The learning process described in Section 4.1 is then repeated using the augmented dataset until convergence. We summarize the full algorithm in Appendix A.
@@ 5.1 Setup
We evaluate our method on four generative question-answering (QA) tasks: three open-domain QA datasets–TruthfulQA (Lin et al., 2022a), TriviaQA (Joshi et al., 2017), and NQ Open (Kwiatkowski et al., 2019); and a domain-specific QA dataset–SciQ (Welbl et al., 2017). For evaluation, 25% of the QA pairs from each dataset are reserved for testing. Consistent with Du et al. (2024), 100 QA pairs are used for validation, while the remaining samples simulate the unlabeled training dataset. We randomly sample $N=32$ pairs from TruthfuQA, and 64 pairs from the other datasets to construct an exemplar set, with $K=128$ used for all experiments. Implementation details are provided in Appendix B.
@@
Evaluation. Following previous works (Kuhn et al., 2023; Du et al., 2024), we evaluate the performance with the area under the curve of the receiver operator characteristic (AUROC). We consider the generation truthful when the similarity score between the generation and the reference answer is larger than a pre-defined threshold (e.g., 0.5). Following Lin et al. (2022a), we utilize BLEURT (Sellam et al., 2020) to measure the similarity. Additionally, we show that our method is robust when evaluated using GPT-4o (Hurst et al., 2024) in Section E.2.
@@ 5.2 Main results
In Table 1, we compare TSV with competitive hallucination detection methods from the literature. TSV demonstrates state-of-the-art performance, significantly outperforming other methods on both the LLaMA-3.1-8b and Qwen-2.5-7b models. We show that unsupervised methods often struggle with inconsistent performance across different models and data distributions as the representations in LLMs are not inherently aligned with the hallucination detection task, making them less reliable for safety-critical applications. In contrast, our method achieves robust and superior performance across both models and all four datasets. In particular, TSV outperforms HaloScope by 13.6% on TruthfulQA with LLaMA-3.1-8b. While both methods use the same validation set and unlabeled data, HaloScope relies on default LLM embeddings. By contrast, our method leverages a small exemplar set and shapes the latent space to better align with the hallucination detection task, enabling significantly improved performance while remaining practical. Our method is also computationally efficient at the inference stage with a complexity of $O(m^{2})$, where $m$ is the number of generated tokens. In contrast, some logit and consistency-based methods require multiple sampling, resulting in a higher complexity of $O(Am^{2})$, where $A$ can be over $10$ in practice. Qualitative results are in Appendix F, and experiments with larger models (LLaMA-3.1-70b & Qwen-2.5-14b) are provided in Section E.1.
@@ 5.3 Ablation studies
Computational efficiency of TSV. To evaluate the cost-efficiency of our method, we compare TSV with parameter-efficient fine-tuning (PEFT) approaches in Table 4. Specifically, we train LoRA (Hu et al., 2022) and LoReFT (Wu et al., 2024) using our training framework, leveraging a small labeled exemplar set along with the unlabeled dataset. Our method achieves superior performance while utilizing 8×8$\times$8 × to 800×800$\times$800 × fewer parameters, demonstrating that TSV can effectively shape representations for the hallucination detection task while significantly reducing computational and annotation costs. Training time is detailed in Appendix H.
@@ 6 Conclusion
Ensuring the reliability of LLM is paramount as they are increasingly integrated into high-stakes applications like healthcare, law, and education. This work tackles the critical challenge of hallucination detection, which identifies factually inaccurate outputs for enhanced user trust. We propose a practical method that minimizes computational and labeling costs while enabling a plug-and-play approach for pre-trained LLMs. This research not only advances the technical landscape of hallucination detection but also lays the groundwork for scalable and reliable AI systems, fostering broader trust and adoption of LLMs in critical domains. Our study does not involve human subjects, complies with all legal and ethical standards, and we do not anticipate any potential harmful consequences resulting from our work. Code is available at: https://github.com/deeplearning-wisc/tsv.
@@ Acknowledgement
We gratefully acknowledge Maxim Khanov, Shawn Im, and Shrey Modi for their valuable comments on the draft. Seongheon Park, Xuefeng Du, Min-Hsuan Yeh and Yixuan Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS2331669, Alfred P. Sloan Fellowship, and Schmidt Sciences Foundation.
@@
Appendix A Algorithms
A.1 Overall training framework
@@ A.2 Sinkhorn algorithm
Appendix B Implementation Details and Hyperparameters
B.1 Implementation details (ours)
Following Kuhn et al. (2023), we generate the most likely answer using beam search with 5 beams. Class prototypes $\boldsymbol{\mu}_{c}$ and TSV $\mathbf{v}$ are randomly initialized, and trained in two stages: 20 epochs using only the exemplar set, followed by an additional 20 epochs after augmentation. Training is performed using the AdamW optimizer (Loshchilov, 2019), with a learning rate of 5e-03 and a batch size of 128. We set steering strength $\lambda$ to 5, the concentration parameter of the vMF distribution $\kappa$ to 10, and the EMA decay rate $\alpha$ to 0.99. The number of iterations in the Sinkhorn algorithm is 3, and the regularization parameter $\epsilon$ is set to 0.05, following Caron et al. (2020). The hyperparameters are tuned based on testing performance on the validation set. The steering location for each model is detailed in Section B.2. For generating responses, we utilize the following input prompt:
Input prompt for generating responses Prompt:
Answer the question concisely:
Q: {question} A:
B.2 Hyperparameters
Table 5: Steering layer index for LLaMA-3.1-8b and Qwen-2.5-7b.
Model Datasets
TruthfulQA TriviaQA SciQ NQ Open
LLaMA-3.1-8b 9 4 8 9
Qwen-2.5-7b 4 6 11 6
Hyperparameters Search space
Steering MHA component {‘mlp’, ‘attn’, ‘res’}
Steering strength ($\lambda$) {0.1, 0.5, 1, 5, 10}
Optimizer {SGD, Adam, AdamW}
Learning rate {1e-04, 2e-04, 5e-04, 1e-03, 2e-03, 5e-03, 1e-02}
Batch size {32, 64, 128}
Initial training epochs ($n_{\text{initial}}$) {5, 10, 20, 40}
Augmented training epochs ($n_{\text{augmented}}$) {5, 10, 20, 40}
EMA decay rate ($\alpha$) {0, 0.5, 0.9, 0.95, 0.99, 1}
Concentration parameter ($\kappa$) {0.1, 1, 5, 10, 100}
B.3 Implementation details (baselines)
Verbalized Prompt:
Q: {question}
Self-evaluation Prompt:
Q: {question}
Appendix C More Details of the Benchmarks
Appendix D Additional Related Works
Hallucination in Large Vision-Language Models (LVLMs). Leveraging the progress in LLMs, Large Vision-Language Models (LVLMs) (Liu et al., 2023; Zhu et al., 2023; Tong et al., 2024; Qiao et al., 2024) have demonstrated strong capabilities in interpreting and reasoning about real-world visual content. Despite these advancements, a fundamental challenge remains in object hallucinations (Rohrbach et al., 2018; Li et al., 2023), where the model incorrectly mentions objects that are not present in the image, often producing outputs that appear plausible but are factually inaccurate.
A growing body of work aims to detect and mitigate object hallucinations in LVLMs by using external models (Liu et al., 2024) or fine-tuning (Sun et al., 2024), but these approaches are often computationally expensive and resource-intensive. Inspired by activation engineering techniques in LLMs, recent approaches (Jiang et al., 2025a, b; Chen et al., 2025; Duan et al., 2025) instead leverage the latent representations within LVLMs to address object hallucinations in a more efficient and interpretable manner. For instance, Nullu (Yang et al., 2025) extracts low-rank subspaces of the differences between truthful and hallucinated features, and further edits the LVLM’s weights to mitigate object hallucinations. VTI (Liu et al., 2025) proposes to steer latent representations during inference to improve the alignment between vision features and textual outputs, thereby reducing object hallucinations.
Appendix E Ablation Studies
E.1 Scalability to larger language models
Table 7: Hallucination detection results on larger LLMs.
Method LLaMA-3.1-70b Qwen-2.5-14b
TruthfulQA SciQ TruthfulQA SciQ
Perplexity 52.3 53.5 64.7 76.0
CCS 63.1 55.5 69.4 80.0
HaloScope 67.4 60.3 74.6 78.9
$\text{SAPLMA}^{\clubsuit}$ 70.2 58.4 83.1 83.8
Ours 76.6 75.0 83.6 89.7
E.2 Evaluation results with GPT-4o
Method LLaMA-3.1-8b Qwen-2.5-7b
TruthfulQA SciQ TruthfulQA SciQ
Perplexity 65.8 41.3 56.3 55.3
CCS 59.5 52.7 58.6 50.3
HaloScope 63.3 66.9 63.4 77.0
$\text{SAPLMA}^{\clubsuit}$ 70.7 72.6 67.7 79.6
TSV (Ours) 75.1 81.6 87.2 85.6
Input prompt for GPT-4o based evaluation Prompt:
E.3 Design choices for the class distribution $\mathbf{w}$
Table 9: Hallucination detection results using different $\mathbf{w}$.
Method LLaMA-3.1-8b Qwen-2.5-7b
TruthfulQA SciQ TruthfulQA SciQ
Uniform 83.2 84.0 87.0 81.2
Estimation 83.7 83.9 87.3 80.6
Oracle 84.3 85.0 87.5 82.6
Ours 84.2 85.8 87.3 82.0
E.4 Results with LLaMA-2-chat-7b
Method TruthfulQA TriviaQA
Perplexity 56.77 72.13
LN-Entropy 61.51 70.91
Semantic Entropy 62.17 73.21
Lexical Similarity 55.69 75.96
EigenScore 51.93 73.98
SelfCKGPT 52.95 73.22
Verbalize 53.04 52.45
Self-evaluation 51.81 55.68
CCS 61.27 60.73
HaloScope 78.64 77.40
TSV (Ours) 80.93 85.20
E.5 Robustness to pseudo-label noise
Table 11: Robustness to pseudo-label noise on LLaMA-3.1-8b.
[…]
We evaluate the impact of pseudo-label noise on our method’s performance. Specifically, we use the same set of selected unlabeled examples from TruthfulQA and systematically introduce noise by flipping some of the correct pseudo-labels. As shown in Table 11, the hallucination detection AUROC on LLaMA-3.1-8b gradually decreases as noise increases, but the performance remains relatively robust, with only a modest drop (from 84.2% to 81.3%) even under 25% label noise. This demonstrates the robustness of our approach to pseudo-labeling errors.
E.6 Class distribution mismatch
Table 12: Ablation study on class distribution mismatch using LLaMA-3.1-8b.
Method TruthfulQA SciQ
Aligned (Ours) 84.2 85.8
Uniform 82.8 82.5
Reversed 82.1 82.7
We analyze the impact of class distribution mismatch between the assumed class distribution (i.e., the exemplar set used for guidance) and the actual class distribution in the unlabeled data. To simulate this, we manually construct exemplar sets under three scenarios: (1) a distribution aligned with the unlabeled generations, (2) a uniform distribution across classes, and (3) a reversed distribution relative to the unlabeled data. As shown in Table 12, our method exhibits a slight performance degradation under mismatched conditions, yet remains competitive across both datasets. This suggests that while alignment between exemplar and target class distributions is beneficial, our method is reasonably robust under class distribution mismatch.
Appendix F Qualitative Results
Appendix G Embedding Norms
Appendix H Compute Resources and Time
H.1 Software and hardware